OpenStudy (anonymous):

Tony has 6 dollars in change in his car. Every time he drives to visit his grandma, he pays $1.25 in tolls. Which graph represents the change Tony has in his car?

OpenStudy (anonymous):

If x = the number of times he drives to his grandma y = -1.25x + 6

OpenStudy (anonymous):

Oh really...thanks I have no idea what to do.

OpenStudy (anonymous):

Can I ask one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which function below has the smallest slope? f(x) f(x) = 2x - 3 g(x) y=2x-5 h(x) x h(x) 1 3 -1 7 3 11 and the options are: f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

@greenleaf800073 ?

OpenStudy (anonymous):

It's function h, look at the number before the 'x', in this case for functions f and g, it's 2 to solve for slope, do (y2-y1)/(x2-x1) y2 is the y coord in the second point, y1 is the y coord in the first point x2 is the x coord in the second point, x1 is the x coord in the firs tpoint In this case, we'll take the points (1, 3) and (-1, 7) (7-3)/(-1-1) = 4/-2 = -2 If you count negative numbers as smaller than positive numbers, then function h has the smallest slope as its slope is -2

OpenStudy (anonymous):

there is an easier way u can but the equation on this form y=mx+c and the slope is m

OpenStudy (anonymous):

No, that's how you do it, and y=mx+c is the equation format

OpenStudy (anonymous):

Oh I see but what if its asking for which function has the lowest y-intercept? it gives: f(x) y=-(1/2)x+2 g(x) Al had two dozen donuts when he walked into the office, but every cubicle he passed, he lost 2 donuts. h(x) h(x) = 6x + 1

OpenStudy (anonymous):

It would be h(x), you see that number after the + sign? that's the y intercept for g(x), he started off with 24 donuts, that 24 donuts would be c in the equation y=mx+c

OpenStudy (anonymous):

Oh ok Thanks ^.^

OpenStudy (anonymous):

What would be the equation for this: Lance gains 1 level each time he plays his favorite video game. He started at level 10.

OpenStudy (anonymous):

It would be y = x+10 x = number of times he plays y = lvl